OpenStudy (anonymous):

What number increased by 1/2 of itself equals 54?

OpenStudy (anonymous):

That should be \(n + \frac{n}{2} = 54\)

OpenStudy (anonymous):

\[\frac{3}{2}n = 54 \Longrightarrow n = 54*\frac{2}{3} = 36\]

hero (hero):

Sorry, I didn't see the "of itself" part

hero (hero):

:/

OpenStudy (anonymous):

let the number be x 1/2 of itself means 1/2 of x \[x + (1\div2)x = 54 \] now LCM \[(2x + x)/.]\div2 = 54\] \[3x \div\] \[3x = 54\times2\] \[3x=108\] \[x=108\div3\] \[x=36\]

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

sorry the third step is incomplete it is \[3x \div2 = 54\]